OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $14,000. Calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $2,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly amount would be lower?

OpenStudy (anonymous):

Payment Option 2 do you need me to show my work??

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

14,000-2,000 (downpayment) = 12,000.....12,000 / 60 = 2,000 a month this is option 1..................... 14,000 / 72 months 194.44 then 2 % of that is 3.88 add 194.44 plus 3.88 = 198.32 way less than the compition of opion 1

OpenStudy (anonymous):

thanks lets give each other a medal!

OpenStudy (anonymous):

i actually just did this for another guy :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

talkin to cutter and eash more

OpenStudy (anonymous):

??

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

eashmore hasn't said a word yet

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but he is typing

OpenStudy (anonymous):

hurry @eashmore

OpenStudy (anonymous):

uf level 5 and ur smarter!! lol

OpenStudy (anonymous):

ya...... what grade or you in

OpenStudy (anonymous):

11

OpenStudy (anonymous):

im in 6th

OpenStudy (anonymous):

and i can answer your question

OpenStudy (anonymous):

im advanced though

OpenStudy (anonymous):

nooooo 11 yrs old

OpenStudy (anonymous):

i do 12th grade math o you can trust me

OpenStudy (anonymous):

i do 9 th grade math

OpenStudy (anonymous):

doing geometry

OpenStudy (anonymous):

let me guess your in 5th grade or six

OpenStudy (anonymous):

im 6th

OpenStudy (anonymous):

Option 1 is pretty easy since there is no interest incurred: \[Monthly~ payment = {(Cost - Down ~payment) \over months}\] Option 2 is a little bit more tricky. \[Monthly~ payment = {(Cost*(1+rate*time)) \over months}\] Remember that time is in years.

OpenStudy (anonymous):

ya so am i

OpenStudy (anonymous):

LETS GIVE MEDALS!!!!!!!!

OpenStudy (anonymous):

i gave you one and you gave me one

OpenStudy (anonymous):

eashmore give me medal and i'll give u one

OpenStudy (anonymous):

gtg ttugl (got to go talk to u guys later)

OpenStudy (anonymous):

lets post that question and give each other medal again